OpenStudy (anonymous):

identify the vertex y = -5x^2 + 10x + 3

OpenStudy (anonymous):

use formula for axis of symmetry x =-b/2a then plug in x to get y and make ordered pair

OpenStudy (anonymous):

a=-5 b=10

OpenStudy (anonymous):

-10/2(-5) = -10/-10= 1

OpenStudy (anonymous):

y=-5(1)^2+10(1)+3 y = -5+10+3 y=5+3 y=8 (1,8) is the vertex